OpenStudy (anonymous):

The sides of a square are two to the power of four ninths inches long. What is the area of the square?

OpenStudy (anonymous):

\[\left(2^{4/9}\right)^2=2^{8/9}=1.85175 \]

OpenStudy (anonymous):

Thankyou!